OpenStudy (anonymous):

can someone explain how to find the limit x-->3 using (9-x^2)/(x-3)?

OpenStudy (anonymous):

Think you have this : 7x = 70 How do you answer ?!:)

OpenStudy (anonymous):

x=10

OpenStudy (anonymous):

Excelent ! Now when we want to find x in your problem can use this way but do you want to know a better way ?!:) :D

OpenStudy (anonymous):

wait if i solve for x is that going to give me the limit as x approaches 3?

OpenStudy (anonymous):

Exacly friend !:) :D

OpenStudy (anonymous):

okay thank you!

OpenStudy (amistre64):

that might be a little misleading ... the limit of a function does not care what the value of the function is

OpenStudy (amistre64):

since at x=3, we get 0/0 the value is undetermined. the limit is defined to be the value that the function approaches, from the left and right. regardless of the actual value of the function

OpenStudy (amistre64):

in this case, we can algebra the top to provide an equivalent setup that cancels out the bottom; so that we do not possess a divide by zero result

OpenStudy (amistre64):

what does the top factor out to be? consider that (x-3) is a factor.

OpenStudy (phi):

as amistre said, a good step is factor the top in \[ \frac{(9-x^2)}{(x-3)} \] notice the top is a "difference of squares", a^2 - b^2 which factors into (a+b)(a-b) if you factor you get \[ \frac{(3+x)(3-x))}{(x-3)} \] notice that you can cancel (x-3) from the top and bottom (as long as x is not *exactly* 3) you get 3+x now let x -->3 what value are you approaching ?

OpenStudy (anonymous):

-6! Sweet thanks so much you guys!!!!

OpenStudy (phi):

I hope you are not saying 3+3 = -6 don't you mean +6 ?